United States currency is printed using intaglio presses thatgenerate a printing pressure of 8.00Image for United States currency is printed using intaglio presses thatgenerate a printing pressure of 8.00 x10^4 lb/in^104 lb/in2. A $20 bill is 6.1in. by 2.6 in. Calculate the magnitude of the force that theprinting press applies to one side of the bill.
Answer #1

Pressure \((P)=8.0 \times 10^{4} l 6(\operatorname{lnch})^{2}\)

Now, the area is given by

$$ \begin{array}{l} A=6.1(\text { inch }) \times 2.6(\text { inch }) \\ o r=15.96(\text { inch })^{2} \end{array} $$

So, magnitude force \((\mathrm{F})=\) A.P

$$ \begin{array}{l} =\left(15.86(\mathrm{inch})^{2}\right) \times\left(8.0 \times 10^{4} 16 .(\mathrm{inch})^{2}\right) \\ F=1.27 \times 10^{4} \mathrm{lb} \\ F=1.27 \times 10^{4} \mathrm{lb} \end{array} $$